# Lost in Overlap: Exploring Watermark Collision in LLMs

### v1

@@ 4.3 Experimental Results
It should also be noted that different watermarking methods behave differently during competition. KGW appears to be less competent than the other two methods. SIR, however, shows significant collisions even in weak paraphraser settings (Column SIR of Tab. 0(a) & 0(c)), while PRW exhibits extreme collisions in strong paraphraser settings (Column PRW of Tab. 0(b) & 0(d)).
Upstream watermarks are prone to be erased by downstream watermarks.
The degradation of the original watermark becomes more pronounced with the increasing strength of the paraphrase watermark, as indicated in Tables 0(b) and 0(d). Regardless of the watermarker’s strength, there is a substantial decrease in the detection accuracy of the paraphrased text $T_{P}$ (observed in the $D_{W}$ columns of Table 1). Upstream watermarks experience heightened vulnerability as the downstream paraphraser strengthens, rendering them nearly indistinguishable from unwatermarked text.
If downstream watermarker has a stronger base model than the upstream watermarker, upstream watermarks can be easily erased.
This observation is supported by the use of LLaMA-2-13B and OPT-1.3B as the base models for the watermarker while maintaining LLaMA-2-13B as the base model for the paraphraser, as illustrated in Fig 3. The findings suggest that watermark collision is inevitable across model sizes. Still, the impact of paraphrase attacks becomes more pronounced when the paraphraser’s base model significantly outperforms the watermarker’s.
Figure 3: TPR for watermark detection with various base models for the watermarker, while employing LLaMA-2-13B as the base model for the paraphraser. FPR is set to 1% and both the watermarkers and paraphrasers are configured with weak settings.
5 Discussion
@@
Since LLM watermarks may be used in a variety of downstream applications, particularly in QA scenarios, collisions between watermarks may cause greater damage to downstream watermarks. Texts with watermarks are more likely to be used as user input or dialogue context. As a result, LLM service providers may have difficulty watermarking their output. This situation may occur more frequently in the future and harm the development of LLM watermarks.
Malicious paraphrase attacks based on watermark collisions.
Previous works Kirchenbauer et al. (2023b, a); Kuditipudi et al. (2023) have introduced several attacks, such as copy-paste attacks and paraphrase attacks, but most have shown their robustness and security against at least some of these attacks. Our study, however, provides a feasible method of constructing effective paraphrase attacks using watermark collisions. The use of paraphrases with strong watermarks could easily wipe out existing watermarks, resulting in greater vulnerability to watermarking techniques. This also presents a significant challenge for watermarking techniques development in the future.
@@
In Table 2, we demonstrate that weak watermarks can still be easily applied to unwatermarked text. It is, however, much more difficult to apply a weak watermark to text that has already been watermarked (Tab. 0(a) & 0(c)). This provides a simple probabilistic method of detecting watermarks without the need to know their details. When it is difficult to add a weak watermark to a paraphrased sentence, it is more likely to have an existing watermark in the original sentence. Thus, watermark collisions could pose a non-negligible threat to the security of watermarking systems.
6 Conclusion
This paper describes the phenomenon of watermark collisions, whereby overlapping watermarks in the same text can reduce the accuracy of both upstream and downstream watermark detection. We conduct our experiments by combining various watermarkers and paraphrasers. In our study, we found that watermark collisions result in overlapping watermarks competing with each other, with one watermark being stronger than the other. We further highlight that watermark collisions may jeopardize the validity and security of watermarks. While we do not intend to resolve this problem of watermark collisions, we hope our work may raise awareness of potential threats to LLM watermarking.
Limitations
Our approach indeed demonstrates the potential collision between existing watermark techniques. Nevertheless, we conduct our experiments only on paraphrasing tasks to simulate watermark collisions. Broader tasks, such as question answering, have not yet been tested.
Furthermore, we select the LLaMA-2-13B as the paraphraser, whereas some other models (e.g. DIPPER) are specially designed for paraphrasing. We believe that the LLaMA has shown its generalization ability in a variety of scenarios. This leads us to believe that it is also capable of paraphrasing.
Also, the maximum number of generated tokens is set to 128. Our research suggests that the observation of watermark collisions would also be evident in a longer context since all logit-based watermarks are susceptible to destructive logit shifts like this. Further experiments are still required.
Acknowledgement
Appendix A Pipeline Setup
@@
  • SIR: We employ the context mode. For weak settings, We set $chunk\_length=10$, $\gamma=0.5$, watermark strenght $\delta=2.0$. For strong settings, we set $chunk\_length=10$, $\gamma=0.5$, watermark strength $\delta=5.0$.
@@
Prompts.
We formulate a prompt tailored for LLaMA-2-13B, enabling it to proficiently paraphrase the given content. The prompt template is shown in Fig. 4.
@@
Assume you are a helpful assistant. Your job is to paraphase the given text.
@@
Figure 4: The paraphrase template for paraphraser.
Appendix B Experimental Results
Table 3 shows the TPR results of detection when utilizing LLaMA-2-13B as the base model of watermarker. Table 4 shows the TPR results of detection when utilizing OPT-1.3B as the base model of watermarker.
@@ Appendix B Experimental Results
Appendix C Scientific Artifacts
The licenses for all the watermarking methods are listed below: KGW (Apache 2.0 Licence), SIR (MIT Licence), PRW (MIT Licence). The licenses for models are listed below: LLaMA-2-13B (LLAMA 2 Community License), OPT-1.3B (OPT LICENSE).

### v2

@@ Introduction
It should also be noted that different watermarking methods behave differently during competition. KGW appears to be less competent than the other two methods. SIR, however, shows significant collisions even in weak paraphraser settings (Column SIR of Tab. 1(a) & 1(c)), while PRW exhibits extreme collisions in strong paraphraser settings (Column PRW of Tab. 1(b) & 1(d)).
However, some anomalies can be observed in back-translation (Table 2), where the TPR increases after a collision. These anomalies occur because certain watermarks may have similar distributions in specific contexts, rendering them ineffective and making attacks on them pointless: When two watermark methods exhibit similar distributions, the likelihood of collision decreases, leading to an increase in the True Positive Rate (TPR). For example, if two KGW-based watermark methods utilize the a similar but not the same red-green list, paraphrasing words into other similar words may accidentally reinforce the watermark within the text, thereby enhancing the TPR. But this is a degraded performance for a certain watermark method: It is a challenge to identify the originating method of the watermark. In practical scenarios, if the origin of the watermark cannot be determined, it essentially means we cannot ascertain which entity embedded the watermark. Consequently, such cases are considered failures of specific watermark methods and do not concern us.
MnF exhibits a similar trend as the watermark collision intensifies, however, it experiences less impact compared to other colliders.
We note that attacks on MnF are less effective because the unmasked words retain most of the context and keep the watermark unaltered. Therefore, the TPR in these cases remains nearly the same. Nonetheless, it is observed that when considering the z-score of $H_{0}$ as stated in (Kirchenbauer et al. 2023a), the influence of watermark collision persists.
In Table 5, we present the z-scores for each corresponding method. Degradation of the z-score is observed after collision attacks compared to those without, indicating the effectiveness of collisions. Besides, the mask rate has a greater impact on the detection process. If the mask rate is low, MnF is less likely to have a significant effect, thus explaining why these attacks have minimal impact on MnF. Furthermore, the SIR method is excluded from MnF experiments, as it prioritizes semantic factors and is less susceptible to MnF attacks.
$W$ $M$varnothing KGWwk PRWwk KGWsg PRWsg
KGWwk 4.14±1.57 3.88±1.51 3.62±1.52 3.52±1.51 3.42±1.56
PRWwk 5.14±1.51 4.89±1.52 4.64±1.59 4.35±1.60 4.44±1.57
KGWsg 6.72±1.95 6.37±1.95 6.09±1.99 5.97±1.87 5.86±1.95
PRWsg 7.82±1.70 7.62±1.74 7.35±1.81 7.08±1.87 7.07±1.82
Table 5: Z-scores of the mask-and-filled text $T_{M}$ with dual watermarks with a mask rate of $0.6$ for masked language modeling tasks. $\varnothing$ means no MnF is applied, i.e., the original detection results of z-scores.
Multi-round Collision
To further enhance the performance of these attacks, multi-round collisions can be applied. We tested the performance of multi-round attacks both with and without watermark collisions, and the results are presented in Fig. 3. In each experiment, we use the same paraphraser to assess the chain effect caused by watermark collisions. For various watermark methods, the TPR of each watermark detection decreases after multi-round collision attacks. As explained in the previous section, applying multi-round collisions causes the word distribution to deviate more significantly from subsequent distributions. The stronger the watermark, the less likely it is for the multi-round watermark to coexist with others.
Possible Application
@@
Since LLM watermarks may be used in a variety of downstream applications, particularly in QA scenarios, collisions between watermarks may cause greater damage to downstream watermarks. Texts with watermarks are more likely to be used as user input or dialogue context. As a result, LLM service providers may have difficulty watermarking their output. This may occur more frequently and harm the development of LLM watermarking techniques.
Malicious attacks based on watermark collisions.
Previous works (Kirchenbauer et al. 2023b, a; Kuditipudi et al. 2023) have introduced several attacks, such as copy-paste attacks and paraphrase attacks, but most have shown their robustness and security against at least some of these attacks. Our study, however, provides a feasible method of constructing effective attacks using watermark collisions. For example, text with a KGWwk watermark can be detected with a 44.8% TPR after a paraphrase attack without a watermark. However, if the paraphrase attack is conducted with another KGWwk, the detection TPR drops to 41.9%. If the attack is done with a KGWsg, the detection TPR further decreases to 28.5%, as presented in Table 1(a) and 1(b). The use of paraphrase, back-translation, and mask-and-fill colliders with strong watermarks could easily wipe out existing watermarks, resulting in greater vulnerability to watermarking.
(a) $P=\text{KGW\textsubscript{{wk}}}$
(b) $P=\text{PRW\textsubscript{{wk}}}$
(c) $P=\text{KGW\textsubscript{{sg}}}$
(d) $P=\text{PRW\textsubscript{{sg}}}$
Figure 3: Multi-round TPR of paraphrased text under a series of paraphrase attacks by the same type of paraphraser with different watermarks. $\varnothing$ represents the original detection TPR before paraphrasing. A sequence of paraphrasers $(P^{(0)},P^{(1)},P^{(2)},\dotsc)$ is applied consecutively to the generated text from the preceding paraphraser.
@@
In Table 1 and 2, we demonstrate that weak watermarks can still be easily applied to unwatermarked text. It is, however, much more difficult to apply a weak watermark to text that has already been watermarked (Tab. 1 & 2). For example, KGW can be applied to plaintext with a success rate of 99.90%, but can be applied to a SIR-watermarked text with a probability not exceeding 25.10%, as presented in Table 1(a) and 1(c). This provides a simple probabilistic method of detecting watermarks without the need to know their details. When it is difficult to add a weak watermark to a paraphrased sentence, it is more likely to have an existing watermark in the original sentence.
Discussion
The LLM watermarking technique is currently undergoing rapid development, with many foundational aspects not yet implemented for practical use regularly. However, our research demonstrates that when watermarks collide, it can significantly hinder the performance of the watermark when applied in real-world situations. A list of predictable risks in practical applications is provided:
  • API Tracing: LLM providers can use watermarking techniques on their LLM API to prevent unauthorized use. However, if the watermarked output of LLMs is sent to other providers with watermarks for further processing, the upstream watermarks will not be effective in tracing the use of the upstream APIs.
  • Black-box Detection: As we discussed in the Experiments Section, the watermark collisions could be used to perform black-box detection of any existing watermarks in any text. Users would experience distrust when they become aware of the presence of watermarks. Hackers may attempt to bypass the watermarks.
Conclusion
In this study, we examine how overlapping watermarks in the same text can decrease the accuracy of both upstream and downstream watermark detection. We propose the use of watermark collisions as an attacking philosophy and therefore emphasize that watermark collisions may compromise the validity and security of all logit-based watermarks. We conduct experiments by integrating various watermarkers and colliders, assessing the text quality before and after collisions, as well as with and without collisions. Our results demonstrate that collision can enhance common attacks and text quality remains consistent when compared to attacks without collisions. While we do not aim to solve the issue of watermark collisions, we hope our work will increase awareness of potential threats to LLM watermarking.
Appendix
Pipeline Setup
@@ Conclusion
[…]
@@
{mdframed}
@@
Assume you are a helpful assistant. Your job is to paraphrase the given text.
@@
Figure 4: The paraphrase prompt template for LLaMA-2 paraphraser.
{mdframed}
<|im_start|>system
You are a helpful assistant. Your job is to paraphrase the given text.<|im_end|>
<|im_start|>user
{INPUT_TEXT}<|im_end|>
<|im_start|>assistant
Figure 5: The paraphrase prompt template for Qwen2 paraphraser.
{mdframed}
[INST] <<SYS>>
Assume you are a helpful assistant.
Your job is to translate the given text from LANGUAGE to LANGUAGE.
<</SYS>>
{INPUT_TEXT} [/INST]
You’re welcome! Here’s a translated version of the original text:
Figure 6: The back-translation prompt template for LLaMA-2 translator.
Prompts.
We formulate a prompt tailored for LLaMA-2-13B, enabling it to proficiently paraphrase the given content. The prompt template is shown in Fig. 4, Fig. 5, and Fig. 6.
Experimental Results
Table 6 and 8 shows the TPR results of detection when utilizing LLaMA-2-13B as the base model of watermarker. Table 7 shows the TPR results of detection when utilizing OPT-1.3B as the base model of watermarker. Table 9 shows the TPR results of detection when utilizing OPT-1.3B as the base model of watermarker.
Collision can be observed across different base models.
This observation is supported by the use of LLaMA-2-13B, Qwen2-7B, and OPT-1.3B as the base models, as illustrated in Fig. 7. The findings suggest that watermark collision is inevitable across different base models, proving its universal applicability as a methodology.
(a) weak $W$, weak $P$
(b) weak $W$, strong $P$
(c) strong $W$, weak $P$
(d) strong $W$, strong $P$
Figure 7: TPR of the paraphrased text $T_{P}$ with different settings and base models as both the watermarker and paraphraser. Blue bars represent TPRs detected by the original watermark detector $D_{W}$, and orange bars are TPRs detected by the collider $D_{P}$. Symbols ,,$\heartsuit$,$\diamondsuit$,$\clubsuit$,$\spadesuit$ , ♢ , ♣ , ♠ denote different watermarker-paraphraser pairs as KGW-KGW, KGW-SIR, SIR-PRW and PRW-KGW, respectively.
@@
Scientific Artifacts
The licenses for all the watermarking methods are listed below: KGW (Apache 2.0 Licence), SIR (MIT Licence), PRW (MIT Licence). The licenses for models are listed below: LLaMA-2-13B (LLAMA 2 Community License), Qwen2-7B-Instruct (Apache 2.0 Licence), OPT-1.3B (OPT LICENSE).
Watermarker KGWwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 99.9 44.8 41.9 9.2 28.5 69.5 34.2 13.7 18.5 58.5 31.4 5.15 3.9 93.5
5% 100 69.7 66.5 19.6 51.2 79.1 59.8 35.8 40.7 82.8 56.8 30.37 14.6 95.56
10% 100 78.4 75.2 30.3 63 85.5 70.3 50.1 55 89.9 68.4 46.92 21.8 96.08
Watermarker KGWsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 100 69.3 67.8 8.1 50 68.9 61.9 13.6 38.8 53 55.3 4.09 7.5 91.7
5% 100 80.7 80.1 25.3 66.8 83 76 40.6 56.7 84.3 70.8 36.09 20.8 95.95
10% 100 86.4 84.1 35.2 74.5 88.2 81.8 58.4 65.9 92.5 76.9 52.25 29.6 96.89
Watermarker PRWwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 95.4 32.9 34.1 7.5 23 68 21.8 12.7 14.3 54.6 25.2 5.73 10.8 92.87
5% 100 59.3 61.1 20.6 46 81 45.4 36 31.1 78.9 49.1 25.86 18.7 95.56
10% 100 78 77.7 32 65.1 86.3 67.3 51.7 50.5 88.4 69.1 44.67 31.3 96.49
Watermarker PRWsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 99.7 63.9 64.8 9.6 50.1 67.9 55.2 10.2 36.6 42.4 50.6 3.56 20.1 90.44
5% 100 80.2 81.1 22.5 71.2 81 76.9 32.2 57.7 75 72.8 27.7 33.6 94.49
10% 100 87 86.8 33.7 80.5 87 83.4 52.8 68.3 89 80.7 46.95 45.1 95.43
Watermarker SIRwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 87.9 5.6 4.75 7.7 4.28 67.4 6.12 11.5 4.4 41.6 - - - -
5% 95.3 18.1 14.26 24.7 16.9 84 22.67 31 17.3 72.6 - - - -
10% 96.6 26.3 21.74 37.7 24.75 89.8 32.4 52.8 25.7 86.9 - - - -
Watermarker SIRsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 94.3 3 1.74 9.4 4.15 74.4 3.82 6.7 2.51 22.5 - - - -
5% 97.8 16.2 14.23 30.4 17.2 86.9 19.98 34.4 15.15 73 - - - -
10% 98.3 24.5 23.75 42.2 27.98 91.2 28.71 51.5 23.87 84.5 - - - -
Table 8: TPR of the back-translation text $T_{P}$ with dual watermarks when utilizing LLaMA-2-13B as the base model of watermarker. FPR is set to $1\%$, $2\%$ $5\%$, respectively.
Watermarker KGWwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 99.9 6 5.4 83.3 1.2 99.8 2.8 55.8 0.3 99.4 4.5 94.15 3.8 99
5% 100 18.8 16.6 93 5.7 99.8 10.1 87.4 2.1 99.9 16.7 97.88 14.7 99.4
10% 100 28.2 25.3 96.5 13.1 99.8 17.9 94.1 4.3 99.9 29.2 98.39 25.9 99.8
Watermarker KGWsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 100 19.2 16.6 85.1 3.6 99.9 12 48 2.1 98.9 14.5 82.09 6.5 98.29
5% 100 35 32.5 93.4 12.3 99.9 25.9 88.2 6.5 100 32.5 97.03 20.3 99.8
10% 100 46.8 45 95.3 20.5 99.9 34.5 94.4 11.5 100 45.3 97.95 31.8 100
Watermarker PRWwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 95.4 8.7 5.8 84.2 1.5 99.8 6.6 46.2 2.8 99.5 5.4 94.47 2.3 99.3
5% 100 20.4 16.4 91.4 4.5 99.9 15.8 81.5 9.2 99.8 15.9 97.69 5.9 99.7
10% 100 39.4 29.4 94.9 8.1 99.9 33.3 91.1 21.5 99.8 31.7 98.69 18.9 99.7
Watermarker PRWsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 99.7 18.7 12.7 85.4 1.4 100 13.1 28.8 6.9 97.2 12.9 86.77 16.3 99.3
5% 100 43.1 33.3 93 7.5 100 37.4 82.1 20.4 99.7 36.6 95.39 39 99.7
10% 100 58.1 48.6 96.6 13.8 100 54 91.5 35 100 55.4 97.7 57.8 99.7
Watermarker SIRwk
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 87.9 3.1 2.86 72 0.8 99.9 3.89 43.1 1.52 98.8 - - - -
5% 95.3 11.2 11.03 86.2 4.81 100 10.84 77.8 6.07 100 - - - -
10% 96.6 18.6 16.96 93.4 10.03 100 16.36 88.4 11.23 100 - - - -
Watermarker SIRsg
Translator No Watermark KGWwk KGWsg PRWwk PRWsg SIRwk SIRsg
FPR Gen TPR $D_{W}$ Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR Old TPR New TPR
1% 94.3 2.5 2.28 77.6 0.81 100 2.93 22.4 2.63 95.4 - - - -
5% 97.8 10.7 12.22 89.4 5.85 100 11.83 75.6 7.48 99.8 - - - -
10% 98.3 18.6 21.43 94.4 12.11 100 20.52 88.2 13.04 100 - - - -
Table 9: TPR of the paraphrased text $T_{P}$ with dual watermarks when utilizing QWEN-7B as the base model of watermarker. FPR is set to $1\%$, $2\%$ $5\%$, respectively.